A commercial diffraction grating has 300 lines per mm.

When a student shines a 470 nm laser through this grating, how many bright spots could be seen on a screen behind the grating?

Express your answer as an integer.
Answer #1

d = width of the slit = 10-3/300 = 3.33 x 10-6 m

= wavelength = 470 nm = 470 x 10-9 m

n = order

N = number of spots

= 90

Using the equation

d Sin = n

(3.33 x 10-6) Sin90 = n (470 x 10-9)

n = 7

Number of spots are given as

N = 2n + 1

N = 2(7) + 1

N = 15